The observatory maintenance team have repaired the network fault and restored the connection to our building.
The fault was caused by a broken Ethernet cable between the switch and the fibre media converter.
Now that the link has been restored the jobs completed over the last two nights will be brought back the the UK server. Any jobs requested since the start of the network failure will be sent to the observatory and enter the job queue there.
The web cams are now active again, and the system will operate tonight with an active RTD feed.